OpenStudy (jamanorthi):

I need help Plz what is 5(m-2)+36=-4

OpenStudy (anonymous):

Destribute the 5 into the parentheses, and then solve for m.

OpenStudy (mertsj):

Use the distributive property. What is 5(m-2) ?

OpenStudy (jamanorthi):

5m-10

OpenStudy (mertsj):

Good

OpenStudy (mertsj):

So now you have 5m-10+36=-4 Add the -10 and 36

OpenStudy (jamanorthi):

26

OpenStudy (mertsj):

So now you have 5m+26=-4 Subtract 26 from both sides.

OpenStudy (jamanorthi):

-21

OpenStudy (mertsj):

Do it again. -4-26 = ???

OpenStudy (jamanorthi):

Never min fight it Thanks

OpenStudy (mertsj):

yw